Solve the given equation.
18c2 + 3c – 10 = 0


Answer by checkley77(12569)

18c2+3c–10=0
(6X+5)(3X-2)=0
6X+5=0
6X=-5
X=-5/6 ANSWER.
3X-2=0
3X=2
X=2/3 ANSWER.